a staircase contains three steps each 10 cm high and 20 cm wide.what should be the minimum horizontal velocity of a ball rolling off the uppermost plane so as to hit directly the lowest plane

Grade:

4 Answers

Monil Shah
11 years ago

The speed should be 10root6. First see max height is 30 cm.now use s=half*g*t(square)

Now divide the distance 60 cm(horizontal range ) by time you get.

The concept is that for max height no initial velocity so get time and for horizontal motion speed is uniform so divide by time to get speed

Pallavi
7 years ago
It is wrong................................?....................................................................
prateek
5 years ago
Let A be the point of origin .Then,
Coordinate of x= 40 and coordinate of y =-20
And angle =0°
Thus using equation for path of trajectory we get ,
Y= tan θ x- gx /2(ucos
prateek
5 years ago
 
Let A be the point of origin .Then,
Coordinate of x= 40 and coordinate of y =-20
And angle =0°
Thus using equation for path of trajectory we get ,
Y= tan θ x- gx /2(ucos theta)2
-20= -1000(40)2/2u2
U= 1000(40)
U=200m/s or 2m/s
Hence the minimum velocity wil be2m/s